#1. Use Thevenins Theorem to find the voltage V, in the circuit below. 3 Ω 2Ω Ο Ο + 50 V 6Ω Σ 0.5υ, & 10 Ω ob
Answer #1

Sola. for ai a giren akti 3 Il 212 (22 +0.5V) Sovet 41-12) + a 10 o.svx b As we have, depenele dent some UTH RTH = Iscapply are KrL hee get in hop 0 -311-6(11-32) +50 = 0 - 3 11 -6 Il +612+50? 911-612 = - 10 Apply Kul in loop @ - 2 I 2 -10 (I2Noth 166.67 V Now short ckt ab Therefore 10 s ceill des couded (52 +0.5 Ux) 로프루 T so ISC 22 -311-6 (Ir-Il) +50 20 911-612 =thee fron, Therinins equiratut clef as RTH + 1 + UTH 18+ Vo = TH=166.620Multimeter-XMM1 XMM1 16.667 A A V dB R1 R3 Set... O! 30 20. XMM2 11 10.5 Mho R2 602 R4 100 V1 -50 V HB Multimeter-XMM2 Х 166.

Note- please find me in comment box if you find any difficulty. Please note that in solution mistakenly I have written Rth=Vth/Rth which is Rth=Vth/Isc. As load resistor is not given, output voltage is nothing but thevenins equivalent voltage across load. Simulation circuit is attached, showing same result. Good luck